Two reactions and their equilibrium constants are given.

A + 2B −⇀ 2C K1 = 2.81

2C −⇀ D K2 = 0.194

Calculate the value of the equilibrium constant for the reaction

D −⇀ A + 2B.

K = ?


1
Expert's answer
2021-06-25T05:41:31-0400

Adding the two equations gives:

A + 2B = 2C K1=2.93

+

2C = D K2=0.258

Result:

A + 2B + 2C 2C + D

K1 + K2

Remove 2С:

A + 2B = D

K1 + K2 = 2.93 + 0.258 = 3.188

Reversed reaction:

D = A + 2B

K3 = 1 / 3.188 = 0.314